Determine for which value of "\alpha" the given series converge absolutely, converge conditionally, or diverges.

  1. "\displaystyle\sum_{k=1}^ ∞ ((-1)^k*
\alpha" ^(k)k)/2
  2. "\displaystyle\sum_{k=1}^ ∞ ((-1)^k*
\alpha" k)/k

Can you kindly explain a little bit so that I can try to solve other problems. Thank you so much.


1
Expert's answer
2021-11-09T10:11:11-0500

use the Leibniz sign.

1. "lim_{k \to \text{\oe}} {\alpha}^{k^k}/2=0"

if "\alpha<1" then the row converges

"lim_{k \to \text{\oe}} {\alpha}^{k^k}/2=\text {\oe}"

If "\alpha>1" then the row diverges.


"lim_{k \to \text{\oe}} |{\alpha}^{k^k}/2|=\text {\oe}"

If "\alpha<0" the series converges conditionally.

"lim_{k \to \text{\oe}} |{\alpha}^{k^k}/2|=0"

If "0< \alpha<1"

the series converges absolutely

2. "lim_{k \to \text{\oe}} {\alpha}^{k}/k=0"

If "\alpha<1"

then the row converges absolutely


"lim_{k \to \text{\oe}} {\alpha}^{k}/k=\text {\oe}"

If "\alpha>1" then the row diverges